# On Malle’s conjecture for the product of symmetric and nilpotent groups

### v2

@@ 2.1. Bound for product of two groups
In this section, we develop upon the results in [Wan21, Section 3] in a framework that is appropriate to our applications. Let $S_{1}$ and $S_{2}$ be multisets with entries in positive integers. We associate the following counting functions with these multi-sets for $i=1,2$
(2.2) $F_{i}(X)\mathrel{\mathop{:}}=\#\{s_{i}\in S_{i}\mathrel{\mathop{:}}s_{i}\leq X\}.$
@@
(2.3) $P_{a,b}(X)\mathrel{\mathop{:}}=\#\{(s_{1},s_{2})\in S_{1}\times S_{2}\mathrel{\mathop{:}}s_{1}^{a}s_{2}^{b}\leq X\}.$
@@
Given $r\in(0,\infty)$ we set $n_{i}(r)$ to be the number of times $r$ occurs in the set $S_{i}$. Then from the definition we deduce that,
(2.6) $\begin{split}P_{a,b}(X)&=\sum_{r_{1}^{a}r_{2}^{b}\leq X}n_{1}(r_{1})n_{2}(r_{2})=\sum_{r^{b}\leq X}n_{2}(r)F_{1}((X/r^{b})^{1/a})\\ &\leq C_{1}X^{1/a}\sum_{r^{b}\leq X}\frac{n_{2}(r)}{r^{b/a}}.\end{split}$
Using Abel’s summation formula and $b-a\cdot\alpha>0$.
[…]
Note that the first term and integral both can be easily bounded above in terms of constants $a,b,\alpha,$ and $\beta$. Substituting the upper bound (2.7) back into (2.6), we obtain the result. ∎
@@
for some $\alpha\in(0,1)$ and $b-a\cdot\alpha>0$ then we have that,
@@
where $c\mathrel{\mathop{:}}=\min\{c_{1}c_{2},c_{1}c_{2}a/(b-a\cdot\alpha)\cdot(1-2^{\alpha-b/a})\}$.
@@
From the definition of $P(X)$ we deduce that,
@@
(2.11) $\begin{split}P_{|G|,n,k}(X)=&\#\{(K,L)\mathrel{\mathop{:}}[K\mathrel{\mathop{:}}k]=n,[L\mathrel{\mathop{:}}k]=|G|,\operatorname{Gal}(\widetilde{K}/k)\simeq S_{n},\operatorname{Gal}(L/k)\simeq G,\\ &K\in\mathcal{P}_{1},L\in\mathcal{P}_{2},\text{ and }\operatorname{inv}_{1}(K)^{|G|}\operatorname{inv}_{2}(L)^{n}\leq X\},\end{split}$
@@
(2.12) $\begin{split}X\ll_{k,S_{n}}\#\{K\mathrel{\mathop{:}}[K\mathrel{\mathop{:}}k]=n,\operatorname{Gal}(\widetilde{K}/k)\simeq S_{n},K\in\mathcal{P}_{1},\text{ and }\operatorname{inv}_{1}(K)\leq X\}\\ X^{\alpha}\ll_{k,G}\#\{L\mathrel{\mathop{:}}[L\mathrel{\mathop{:}}k]=|G|,\operatorname{Gal}(L/k)\simeq G,L\in\mathcal{P}_{2},\text{ and }\operatorname{inv}_{2}(L)\leq X\}.\end{split}$
@@
    the bounds (2.10) and (2.12) are satisfied for $\alpha\mathrel{\mathop{:}}=a(G)$,
@@
The result follows from (2.10) and (2.12) and Propositions 3.5 and 3.5. ∎
@@ 3. Local uniformity estimates
In this section, we state local uniformity estimates for degree $n$ extensions of $k$ with Galois group $S_{n}$ given in [Wan21, Section 4] for $n=3,4,$ and, $5$. We refer to an extension $K/k$ as a degree $n$, $S_{n}$-extension if $[K\mathrel{\mathop{:}}k]=n$ and $\operatorname{Gal}(\widetilde{K}/k)\simeq S_{n}$. Furthermore, we use the parameterization for nilpotent extensions developed in [KP23] to prove a similar estimate for $G$-extensions of $K$, where $G$ is any finite nilpotent group. We refer to an extension $L/k$ as a $G$-extension if $L$ is Galois over $k$ and $\operatorname{Gal}(L/k)\simeq G$. Let $\mathfrak{p}$ be a finite prime of $k$, and $K/k$ be an $S_{4}$-extension then, we say that $\mathfrak{p}$ is overramified in $K/k$ if $\mathfrak{p}$ factors into primes in $K/k$ as $\mathfrak{B}^{4},\mathfrak{B}^{2},$ or $\mathfrak{B}_{1}^{2}\mathfrak{B}_{2}^{2}$ for finite $p$. For an archimedean $\mathfrak{p}$, it must factor into a product of two ramified places.
@@ 3.1. A parameterization for nilpotent extensions
We begin by considering the case when $G$ is an $\ell$-group, i.e., $|G|$ is a power of $\ell$. Let $S_{\infty}\subset\Omega_{k}$ be the archimedian primes of $k$. We view $k$ is a subfield of $\bar{\mathbb{Q}}$ and set $\operatorname{G}_{k}$ to denote the Galois group $\operatorname{Gal}(\bar{\mathbb{Q}}/k)$. Likewise, for each prime $\mathfrak{q}\in\Omega_{k}$, denote by $k_{\mathfrak{q}}$ the completion of $K$ at $q$ and set $\operatorname{G}_{k_{\mathfrak{q}}}\mathrel{\mathop{:}}=\operatorname{Gal}(\bar{\mathbb{Q}}_{\mathfrak{p}}/k_{\mathfrak{q}})$. For each prime $q\in\Omega_{k}$, choose an embedding $\iota_{\mathfrak{q}}\mathrel{\mathop{:}}\bar{k}\hookrightarrow\bar{k}_{\mathfrak{q}}$. The choice of embedding induces an inclusion of $\iota_{\mathfrak{q}}^{\ast}\mathrel{\mathop{:}}\operatorname{G}_{k_{\mathfrak{q}}}\hookrightarrow\operatorname{G}_{k}$. Denote by $\operatorname{I}_{q}\subset\operatorname{G}_{k_{\mathfrak{q}}}$ the inertia group at $q$. We set $\mathbb{F}_{\ell}$ to be the field with $\ell$-elements, and set $H^{1}_{\operatorname{nr}}(\operatorname{G}_{k_{\mathfrak{q}}},\mathbb{F}_{\ell})$ is the image of the inflation map
@@
$\Phi(\ell,S)\mathrel{\mathop{:}}H^{1}(\operatorname{G}_{k},\mathbb{F}_{\ell})\rightarrow\bigoplus_{q\in\Omega_{k}\setminus S}\frac{H^{1}(\operatorname{G}_{k_{\mathfrak{q}}},\mathbb{F}_{\ell})}{H^{1}_{\operatorname{nr}}(\operatorname{G}_{k_{\mathfrak{q}}},\mathbb{F}_{\ell})}.$
@@
Every prime $q\in\Omega_{k}\setminus\left(S_{\operatorname{clean}}(\ell)\cup\widetilde{\Omega}_{k}(\ell)\right)$ is unramified in all finite $\ell$-extensions of $K$ (cf. [KP23][Proposition 2.3]). For $\mathfrak{q}\in\widetilde{\Omega}_{k}(\ell)$ there exists a character $\chi_{\mathfrak{q}}\in H^{1}(\operatorname{G}_{K},\mathbb{F}_{\ell})$ such that $\Phi(\ell,S_{\operatorname{clean}}(\ell))$ has nontrivial coordinate precisely at $\mathfrak{q}$ and at no other prime in $\widetilde{\Omega}_{k}(\ell)$. The set of characters $\chi_{\mathfrak{q}}$ (as $\mathfrak{q}$ ranges over $\widetilde{\Omega}_{k}(\ell)$) is a linearly independent set. There is a positive integer $t$ and a basis
$J\mathrel{\mathop{:}}=\{\chi_{i}\mid i=1,\dots,t\}$
@@
Let $[0,1]^{t}$ denote the set of all vectors $\vec{v}=(v_{1},\dots,v_{t})$ whose entries are either 00 or $1$. Given $T\in[0,1]^{t}$, set $\pi_{j}(T)$ to denote its $j$-th coordinate. Set $\mathcal{S}_{\ell}\mathrel{\mathop{:}}=[0,1]^{[t]}\times\mathcal{S}_{\ell}^{\prime}$, where $\mathcal{S}_{\ell}^{\prime}$ is the set consisting of squarefree integral ideals of $\mathcal{O}_{k}$ supported entirely on $\widetilde{\Omega_{k}}(\ell)$. Denote by $\operatorname{Prim}\left(\mathcal{S}_{\ell}^{G^{\ast}}\right)$ the set of vectors supported on $G^{\ast}$ having pairwise coprime coordinates. Here, two pairs $(T,\mathfrak{b})$ and $(T^{\prime},\mathfrak{b}^{\prime})$ are coprime if $\mathfrak{b}$ and $\mathfrak{b}^{\prime}$ are coprime and there does not exist a $j\in[t]$, such that $\pi_{j}(T)=\pi_{j}(T^{\prime})=1$. Elements of $\operatorname{Prim}\left(\mathcal{S}_{\ell}^{G^{\ast}}\right)$ are vectors $(v_{g})_{g\in G^{\ast}}$ where $v_{g}=(v_{g}{(1)},v_{g}{(2)})$. Let $\operatorname{Epi}_{\operatorname{top-gr}}\left(\mathcal{G}_{k}^{\operatorname{pro}-\ell};G\right)$ denote the set of surjective homomorphisms $\mathcal{G}_{k}^{\operatorname{pro}-\ell}\twoheadrightarrow G$.
@@
$P_{G}\mathrel{\mathop{:}}\operatorname{Prim}\left(\mathcal{S}_{\ell}^{G^{\ast}}\right)\twoheadrightarrow\operatorname{Epi}_{\operatorname{top-gr}}\left(\mathcal{G}_{k}^{\operatorname{pro}-\ell};G\right)\cup\{\cdot\}.$
@@
The pre-image of $\operatorname{Epi}_{\operatorname{top-gr}}\left(\mathcal{G}_{k}^{\operatorname{pro}-\ell};G\right)$ is denoted by $\operatorname{Prim}\left(\mathcal{S}_{\ell}^{G^{\ast}}\right)(\operatorname{solv.})$. For an epimorphism $\psi\mathrel{\mathop{:}}\mathcal{G}_{k}^{\operatorname{pro}-\ell}\twoheadrightarrow G$, set $\operatorname{Disc}(\psi)$ to denote the relative discriminant of the corresponding extension of $k$ that is fixed by the kernel of $\psi$. Given an ideal $\mathfrak{b}$, denote by $\operatorname{free}_{S_{\operatorname{clean}}(\ell)}(\mathfrak{b})$ to be the largest ideal that divides $\mathfrak{b}$ and is supported outside $S_{\operatorname{clean}}(\ell)$.
@@
$P_{G}\mathrel{\mathop{:}}\prod_{j\in[c]}\operatorname{Prim}\left(\mathcal{S}_{\ell_{j}}^{G(\ell_{j})-\operatorname{id}}\right)\twoheadrightarrow\operatorname{Epi}_{\operatorname{top}-\operatorname{gr}}\left(\operatorname{G}_{k},G\right)\cup\{\cdot\}.$
@@
$\mathcal{S}\mathrel{\mathop{:}}=\{0,1\}^{[t_{1}+\dots+t_{c}]}\times\mathcal{S}^{\prime},$
@@
  • writing $\Pi_{i}\mathrel{\mathop{:}}[\sum t_{i}]\rightarrow[t_{i}]$ for the natural projection map, then, $\Pi_{i}(v_{g}(1))$ are pairwise coprime,
@@
$P_{G}\mathrel{\mathop{:}}\operatorname{Prim}\left(\mathcal{S}^{G^{\ast}}\right)\twoheadrightarrow\operatorname{Epi}_{\operatorname{top}-\operatorname{gr}}\left(\operatorname{G}_{k},G\right)\cup\{\cdot\}$
@@
$\operatorname{Prim}\left(\mathcal{S}^{G^{\ast}}\right)(\operatorname{solv.})\mathrel{\mathop{:}}=P_{G}^{-1}\left(\operatorname{Epi}_{\operatorname{top}-\operatorname{gr}}\left(\operatorname{G}_{k},G\right)\right).$
@@
$A_{z}(x)\mathrel{\mathop{:}}=\sum_{\begin{subarray}{c}|I|\leq x,\\ \mathfrak{p}|I\Rightarrow\operatorname{Frob}_{\mathfrak{p}}\in\mathfrak{S}\text{ and }\mathfrak{p}\notin\mathcal{P}\end{subarray}}\mu^{2}(I)z^{\omega_{\mathfrak{S}}(I)},$
@@ 3.2. Uniformity estimate for nilpotent groups
Let $G$ be a non-trivial finite nilpotent group. Given an integral ideal $\mathfrak{q}$ of $\mathcal{O}_{k}$, set $N_{q}(G;X)$ to be the number of Galois extensions $L/k$ with $\operatorname{Gal}(L/k)=G$, $\mathrm{Disc}(L)\leq X$, and such that $\mathfrak{q}$ divides $\mathrm{disc}(L/k)$. We recall that $\ell_{G}$ denotes the smallest prime divisor of $|G|$ and $I(G)$ is the subset of elements of $G$ with order $\ell_{G}$. Set $H(G)\mathrel{\mathop{:}}=I(G)\cup\{\operatorname{id}\}$. We establish an extension a uniformity estimate for nilpotent groups. This generalizes [Wan21][Theorem 4.12]. Recall that
$i(G,k)=\frac{\#I(G)}{[k(\mu_{\ell_{G}})\mathrel{\mathop{:}}k]}.$
@@
Without loss of generality, we may assume that $\mathfrak{q}$ is not divisible by any prime in $S$. For $g\in G$, set $e_{g}\mathrel{\mathop{:}}=|G|\left(1-\frac{1}{\#\langle g\rangle}\right)$. Observe that for $g\in I(G)$, we have that $e_{g}=a(G)^{-1}$. It follows from the parameterization in the previous subsection and Proposition 3.4 that it suffices to bound the following
(3.3) #$\#\left\{v=(v_{g}(1),v_{g}(2))\in\operatorname{Prim}\left(\mathcal{S}^{G^{\ast}}\right)\mathrel{\mathop{:}}\mathfrak{q}\mid\Delta_{v}\text{ and }|\Delta_{v}|\leq X\right\},$
@@
$\Delta_{v}\mathrel{\mathop{:}}=\prod_{g\in G-\{\operatorname{id}\}}v_{g}(2)^{e_{g}}.$
@@
$\begin{split}&v^{\prime}\mathrel{\mathop{:}}=\pi_{G-H(G)}(v)\mathrel{\mathop{:}}=\prod_{g\in G-H(G)}(v_{g}(1),v_{g}(2));\\ &v^{\prime\prime}\mathrel{\mathop{:}}=\pi_{I(G)}(v)\mathrel{\mathop{:}}=\prod_{h\in I(G)}(v_{h}(1),v_{h}(2)).\end{split}$
The elements $v^{\prime}$ and $v^{\prime\prime}$ belong to $\operatorname{Prim}\left(\mathcal{S}^{G-H(G)}\right)$ and $\operatorname{Prim}\left(\mathcal{S}^{I(G)}\right)$ respectively. We set $N_{\mathfrak{q}_{2}}(v^{\prime},X)$ to denote the number of $v^{\prime\prime}=\prod_{h\in I(G)}(v_{1}(h),v_{2}(h))_{h\in I(G)}\in\operatorname{Prim}\left(\mathcal{S}^{I(G)}\right)$ such that
@@
Set $\widetilde{\mathfrak{q}}_{2}\mathrel{\mathop{:}}=\prod_{\mathfrak{p}\mid\mathfrak{q}_{2}}\mathfrak{p}$ to be the squarefree part of $\mathfrak{q}_{2}$ and write $\widetilde{\mathfrak{q}}_{2}I_{2}=\prod_{h\in I(G)}v_{h}(2)$. For $h\in I(G)$, recall that the ideals $v_{h}(2)$ are all squarefree and mutually coprime. Furthermore, they are supported on $\widetilde{\Omega}_{K}(\ell_{G})$. Therefore, $\widetilde{\mathfrak{q}}_{2}I_{2}$ is squarefree and supported in $\widetilde{\Omega}_{K}(\ell_{G})$. It is easy to see that
@@
Set $L$ to be the cyclotomic extension $k(\mu_{\ell_{G}})$ and $\mathfrak{S}=\{1\}\subset\operatorname{Gal}(L/k)$. Note that an ideal $I$ is supported on $\widetilde{\Omega}_{K}(\ell_{G})$ if and only if all prime ideals $\mathfrak{p}|I$ split in $L$. Setting $z\mathrel{\mathop{:}}=\#I(G)$, rewrite (3.5) as
@@
$Y\mathrel{\mathop{:}}=\left(\frac{X^{a(G)}}{|\mathfrak{q}_{2}|^{a(G)}|\prod_{g\in G-H(G)}v_{g}(2)^{e_{g}a(G)}|}\right).$
@@
$N_{\mathfrak{q}_{2}}(v^{\prime},X)\ll_{G,k,\epsilon}Y(\log Y)^{\frac{z\#\mathfrak{S}}{[k(\mu_{\ell_{G}})\mathrel{\mathop{:}}k]}-1}.$
@@
$i(G,k)=\frac{\#I(G)}{[k(\mu_{\ell_{G}})\mathrel{\mathop{:}}k]}=\frac{z\#\mathfrak{S}}{[k(\mu_{\ell_{G}})\mathrel{\mathop{:}}k]},$
@@
Observe that $e_{g}a(G)>1$ for $g\in G-H(G)$. Setting $t\mathrel{\mathop{:}}=|G|-\#H(G)$, enumerate the elements of $G-H(G)$ as $\{g_{1},\dots,g_{t}\}$. Setting, $e_{i}\mathrel{\mathop{:}}=e(g_{i})$, we find that

### v3

@@ 2.1. Bound for product of two groups
In this section, we develop upon the results in [Wan21, Section 3] in a framework that is appropriate to our applications. Let $S_{1}$ and $S_{2}$ be multisets with entries in positive integers. Furthermore, assume that $S_{2}$ contains some integer other than $1$. We associate the following counting functions with these multi-sets for $i=1,2$
(2.7) $F_{i}(X):=\#\{s_{i}\in S_{i}:s_{i}\leq X\}.$
@@
(2.8) $P_{a,b}(X):=\#\{(s_{1},s_{2})\in S_{1}\times S_{2}:s_{1}^{a}s_{2}^{b}\leq X\}.$
@@
The result follows from [Wan21, Lemma 3.2], setting $n_{1}:=1/a$ and $n_{2}:=\alpha/b$ (in accordance with the notation in loc. cit.)). We note that the condition $b-a\alpha>0$ translates to $n_{1}>n_{2}$, which is a necessary condition for the Lemma 3.2 of loc. cit. to apply. ∎
@@
for some $\alpha\in(0,1)$ and $b-a\cdot\alpha>0$ then, for all large enough values of $X$, we have that
@@
where $c:=\min\{c_{1}c_{2},c_{1}c_{2}a/(b-a\cdot\alpha)\cdot(1-2^{\alpha-b/a})\}$.
@@
The proof of our result adapts the first part of [Wan21, Lemma 3.2]. Given $r\in(0,\infty)$, we set $n_{i}(r)$ to be the number of times $r$ occurs in the set $S_{i}$. From the definition of $P_{a,b}(X)$, we deduce that
@@
(2.14) $\begin{split}P_{|G|,n,k}(X)=&\#\{(K,L):[K:k]=n,[L:k]=|G|,\operatorname{Gal}(\widetilde{K}/k)\simeq S_{n},\operatorname{Gal}(L/k)\simeq G,\\ &K\in\mathcal{P}_{1},L\in\mathcal{P}_{2},\text{ and }\operatorname{inv}_{1}(K)^{|G|}\operatorname{inv}_{2}(L)^{n}\leq X\},\end{split}$
@@
(2.15) $\begin{split}X\ll_{k,S_{n}}\#\{K:[K:k]=n,\operatorname{Gal}(\widetilde{K}/k)\simeq S_{n},K\in\mathcal{P}_{1},\text{ and }\operatorname{inv}_{1}(K)\leq X\}\\ X^{\alpha}\ll_{k,G}\#\{L:[L:k]=|G|,\operatorname{Gal}(L/k)\simeq G,L\in\mathcal{P}_{2},\text{ and }\operatorname{inv}_{2}(L)\leq X\}.\end{split}$
@@
    the bounds (2.13) and (2.15) are satisfied for $\alpha:=a(G)$,
@@
The result follows from (2.13) and (2.15) and Propositions 2.6 and 2.7. ∎
@@ 3. Local uniformity estimates
[…]
@@ 3.1. A parameterization for nilpotent extensions
We begin by considering the case when $G$ is an $\ell$-group, i.e., $|G|$ is a power of $\ell$. Let $S_{\infty}\subset\Omega_{k}$ be the archimedian primes of $k$. We view $k$ is a subfield of $\bar{\mathbb{Q}}$ and set $\operatorname{G}_{k}$ to denote the Galois group $\operatorname{Gal}(\bar{\mathbb{Q}}/k)$. Likewise, for each prime $\mathfrak{q}\in\Omega_{k}$, denote by $k_{\mathfrak{q}}$ the completion of $K$ at $q$ and set $\operatorname{G}_{k_{\mathfrak{q}}}:=\operatorname{Gal}(\bar{\mathbb{Q}}_{\mathfrak{p}}/k_{\mathfrak{q}})$. For each prime $q\in\Omega_{k}$, choose an embedding $\iota_{\mathfrak{q}}:\bar{k}\hookrightarrow\bar{k}_{\mathfrak{q}}$. The choice of embedding induces an inclusion of $\iota_{\mathfrak{q}}^{\ast}:\operatorname{G}_{k_{\mathfrak{q}}}\hookrightarrow\operatorname{G}_{k}$. Denote by $\operatorname{I}_{q}\subset\operatorname{G}_{k_{\mathfrak{q}}}$ the inertia group at $q$. We set $\mathbb{F}_{\ell}$ to be the field with $\ell$-elements, and set $H^{1}_{\operatorname{nr}}(\operatorname{G}_{k_{\mathfrak{q}}},\mathbb{F}_{\ell})$ is the image of the inflation map
@@
$\Phi(\ell,S):H^{1}(\operatorname{G}_{k},\mathbb{F}_{\ell})\rightarrow\bigoplus_{q\in\Omega_{k}\setminus S}\frac{H^{1}(\operatorname{G}_{k_{\mathfrak{q}}},\mathbb{F}_{\ell})}{H^{1}_{\operatorname{nr}}(\operatorname{G}_{k_{\mathfrak{q}}},\mathbb{F}_{\ell})}.$
@@
Every prime $\mathfrak{q}\in\Omega_{k}\setminus\left(S_{\operatorname{clean}}(\ell)\cup\widetilde{\Omega}_{k}(\ell)\right)$ is unramified in all finite $\ell$-extensions of $K$ (cf. [KP23][Proposition 2.3]). For $\mathfrak{q}\in\widetilde{\Omega}_{k}(\ell)$ there exists a character $\chi_{\mathfrak{q}}\in H^{1}(\operatorname{G}_{K},\mathbb{F}_{\ell})$ such that $\Phi(\ell,S_{\operatorname{clean}}(\ell))$ has nontrivial coordinate precisely at $\mathfrak{q}$ and at no other prime in $\widetilde{\Omega}_{k}(\ell)$, see [KP23, p. 317, line -9] for details. The set of characters $\chi_{\mathfrak{q}}$ (as $\mathfrak{q}$ ranges over $\widetilde{\Omega}_{k}(\ell)$) is a linearly independent set. There is a positive integer $t$ and a basis
$J:=\{\chi_{i}\mid i=1,\dots,t\}$
@@
Let $[0,1]^{t}$ denote the set of all vectors $\vec{v}=(v_{1},\dots,v_{t})$ whose entries are either 00 or $1$. Given $T\in[0,1]^{t}$, set $\pi_{j}(T)$ to denote its $j$-th coordinate. Set $\mathcal{S}_{\ell}:=[0,1]^{[t]}\times\mathcal{S}_{\ell}^{\prime}$, where $\mathcal{S}_{\ell}^{\prime}$ is the set consisting of squarefree integral ideals of $\mathcal{O}_{k}$ supported entirely on $\widetilde{\Omega_{k}}(\ell)$. Recall that $G^{\ast}:=G\backslash\{1_{G}\}$. Two pairs $(T,\mathfrak{b})$ and $(T^{\prime},\mathfrak{b}^{\prime})$ in $\mathcal{S}_{\ell}$ are said to be coprime if $\mathfrak{b}$ and $\mathfrak{b}^{\prime}$ are coprime and there does not exist a $j\in[t]$, such that $\pi_{j}(T)=\pi_{j}(T^{\prime})=1$. Consider the vectors $(v_{g})_{g\in G^{\ast}}$, indexed by $g\in G^{\ast}$, having entries $v_{g}\in\mathcal{S}_{\ell}$. Denote by $\operatorname{Prim}\left(\mathcal{S}_{\ell}^{G^{\ast}}\right)$ the subset of such vectors whose coordinates are pairwise coprime. Elements of $\operatorname{Prim}\left(\mathcal{S}_{\ell}^{G^{\ast}}\right)$ shall be represented as $(v_{g})_{g\in G^{\ast}}$, where $v_{g}=(v_{g}{(1)},v_{g}{(2)})$ with $v_{g}{(1)}\in[0,1]^{t}$ and $v_{g}{(2)}\in\mathcal{S}_{\ell}^{\prime}$. Let $\operatorname{Epi}_{\operatorname{top-gr}}\left(\mathcal{G}_{k}^{\operatorname{pro}-\ell};G\right)$ denote the set of surjective homomorphisms $\mathcal{G}_{k}^{\operatorname{pro}-\ell}\twoheadrightarrow G$.
@@
$P_{G}:\operatorname{Prim}\left(\mathcal{S}_{\ell}^{G^{\ast}}\right)\twoheadrightarrow\operatorname{Epi}_{\operatorname{top-gr}}\left(\mathcal{G}_{k}^{\operatorname{pro}-\ell};G\right)\cup\{\cdot\}.$
@@
The pre-image of $\operatorname{Epi}_{\operatorname{top-gr}}\left(\mathcal{G}_{k}^{\operatorname{pro}-\ell};G\right)$ is denoted by $\operatorname{Prim}\left(\mathcal{S}_{\ell}^{G^{\ast}}\right)(\operatorname{solv.})$. For an epimorphism $\psi:\mathcal{G}_{k}^{\operatorname{pro}-\ell}\twoheadrightarrow G$, set $\operatorname{Disc}(\psi)$ to denote the relative discriminant of the corresponding extension of $k$ that is fixed by the kernel of $\psi$. Given an ideal $\mathfrak{b}$, denote by $\operatorname{free}_{S_{\operatorname{clean}}(\ell)}(\mathfrak{b})$ to be the largest ideal that divides $\mathfrak{b}$ and is supported outside $S_{\operatorname{clean}}(\ell)$.
@@
$P_{G}:\prod_{j\in[c]}\operatorname{Prim}\left(\mathcal{S}_{\ell_{j}}^{G(\ell_{j})-\operatorname{id}}\right)\twoheadrightarrow\operatorname{Epi}_{\operatorname{top}-\operatorname{gr}}\left(\operatorname{G}_{k},G\right)\cup\{\cdot\}.$
@@
$\mathcal{S}:=\{0,1\}^{[t_{1}+\dots+t_{c}]}\times\mathcal{S}^{\prime},$
@@
  • writing $\Pi_{i}:[\sum t_{i}]\rightarrow[t_{i}]$ for the natural projection map, then, $\Pi_{i}(v_{g}(1))$ are pairwise coprime,
@@
$P_{G}:\operatorname{Prim}\left(\mathcal{S}^{G^{\ast}}\right)\twoheadrightarrow\operatorname{Epi}_{\operatorname{top}-\operatorname{gr}}\left(\operatorname{G}_{k},G\right)\cup\{\cdot\}$
@@
$\operatorname{Prim}\left(\mathcal{S}^{G^{\ast}}\right)(\operatorname{solv.}):=P_{G}^{-1}\left(\operatorname{Epi}_{\operatorname{top}-\operatorname{gr}}\left(\operatorname{G}_{k},G\right)\right).$
@@
$A_{z}(x):=\sum_{\begin{subarray}{c}|I|\leq x,\\ \mathfrak{p}|I\Rightarrow\operatorname{Frob}_{\mathfrak{p}}\in\mathfrak{S}\text{ and }\mathfrak{p}\notin\mathcal{P}\end{subarray}}\mu^{2}(I)z^{\omega_{\mathfrak{S}}(I)},$
@@ 3.2. Uniformity estimate for nilpotent groups
Let $G$ be a non-trivial finite nilpotent group. Given an integral ideal $\mathfrak{q}$ of $\mathcal{O}_{k}$, set $N_{q}(G;X)$ to be the number of Galois extensions $L/k$ with $\operatorname{Gal}(L/k)=G$, $\mathrm{Disc}(L)\leq X$, and such that $\mathfrak{q}$ divides $\mathrm{disc}(L/k)$. We recall that $\ell_{G}$ denotes the smallest prime divisor of $|G|$ and $I(G)$ is the subset of elements of $G$ with order $\ell_{G}$. Set $H(G):=I(G)\cup\{\operatorname{id}\}$. We establish an extension a uniformity estimate for nilpotent groups. This generalizes [Wan21][Theorem 4.12]. Recall from [KP23] that
$i(G,k)=\frac{\#I(G)}{[k(\mu_{\ell_{G}}):k]}.$
@@
Without loss of generality, we may assume that $\mathfrak{q}$ is not divisible by any prime in $S$. For $g\in G$, set $e_{g}:=|G|\left(1-\frac{1}{\#\langle g\rangle}\right)$. Observe that for $g\in I(G)$, we have that $e_{g}=a(G)^{-1}$. It follows from the parameterization in the previous subsection and Proposition 3.4 that it suffices to bound the following
(3.3) #$\#\left\{v=(v_{g}(1),v_{g}(2))\in\operatorname{Prim}\left(\mathcal{S}^{G^{\ast}}\right):\mathfrak{q}\mid\Delta_{v}\text{ and }|\Delta_{v}|\leq X\right\},$
@@
$\Delta_{v}:=\prod_{g\in G-\{\operatorname{id}\}}v_{g}(2)^{e_{g}}.$
@@
$\begin{split}&v^{\prime}:=\pi_{G-H(G)}(v):=\prod_{g\in G-H(G)}(v_{g}(1),v_{g}(2));\\ &v^{\prime\prime}:=\pi_{I(G)}(v):=\prod_{h\in I(G)}(v_{h}(1),v_{h}(2)).\end{split}$
The elements $v^{\prime}$ and $v^{\prime\prime}$ belong to $\operatorname{Prim}\left(\mathcal{S}^{G-H(G)}\right)$ and $\operatorname{Prim}\left(\mathcal{S}^{I(G)}\right)$ respectively. We set $N_{\mathfrak{q}_{2}}(v^{\prime},X)$ to denote the number of $v^{\prime\prime}=\prod_{h\in I(G)}(v_{h}(1),v_{h}(2))\in\operatorname{Prim}\left(\mathcal{S}^{I(G)}\right)$ such that
@@
Set $\widetilde{\mathfrak{q}}_{2}:=\prod_{\mathfrak{p}\mid\mathfrak{q}_{2}}\mathfrak{p}$ to be the squarefree part of $\mathfrak{q}_{2}$ and write $\widetilde{\mathfrak{q}}_{2}I_{2}=\prod_{h\in I(G)}v_{h}(2)$. For $h\in I(G)$, recall that the ideals $v_{h}(2)$ are all squarefree and mutually coprime. Furthermore, they are supported on $\widetilde{\Omega}_{K}(\ell_{G})$. Therefore, $\widetilde{\mathfrak{q}}_{2}I_{2}$ is squarefree and supported in $\widetilde{\Omega}_{K}(\ell_{G})$. It is easy to see that
@@
Set $L$ to be the cyclotomic extension $k(\mu_{\ell_{G}})$ and $\mathfrak{S}=\{1\}\subset\operatorname{Gal}(L/k)$. Note that an ideal $I$ is supported on $\widetilde{\Omega}_{K}(\ell_{G})$ if and only if all prime ideals $\mathfrak{p}|I$ split in $L$. Setting $z:=\#I(G)$, rewrite (3.5) as
@@
$Y:=\left(\frac{X^{a(G)}}{|\mathfrak{q}_{2}|^{a(G)}|\prod_{g\in G-H(G)}v_{g}(2)^{e_{g}a(G)}|}\right).$
@@
$N_{\mathfrak{q}_{2}}(v^{\prime},X)\ll_{G,k,\epsilon}Y(\log Y)^{\frac{z\#\mathfrak{S}}{[k(\mu_{\ell_{G}}):k]}-1}.$
@@
$i(G,k)=\frac{\#I(G)}{[k(\mu_{\ell_{G}}):k]}=\frac{z\#\mathfrak{S}}{[k(\mu_{\ell_{G}}):k]},$
@@
Observe that $e_{g}a(G)>1$ for $g\in G-H(G)$. Setting $t:=|G|-\#H(G)$, enumerate the elements of $G-H(G)$ as $\{g_{1},\dots,g_{t}\}$. Setting, $e_{i}:=e(g_{i})$, we find that